# A functional central limit theorem for weighted occupancy processes of the Karlin model

### v2

@@ 1. Introduction
The Karlin model is an infinite urn scheme where the sampling frequencies $p_{j},j\in\mathbb{N}$ decay at a polynomial rate $p_{j}\sim Cj^{-1/\alpha}$ (possibly with a multiplicative slowly varying function) for some parameter $\alpha\in(0,1)$. The infinite urn scheme has been a classical model in probability theory, dating back at least to the 1960s in the works of Bahadur, [3], Darling, [12] and Karlin, [25]. Among many statistics of interest, Karlin established central limit theorems for the so-called occupancy and odd-occupancy processes when the sampling frequencies decay at a polynomial rate. If we let $D_{n,j}$ denote the number of urns with exactly $j$ balls after the first $n$ rounds of sampling, then these are the counting processes
@@
In this paper, we investigated the weighted sum of occupancy counts $(D_{n,j})_{j\in\mathbb{N}}$ with prescribed weights $\vec{a}=(a_{j})_{j\in\mathbb{N}}$ defined as
@@
The statistics in this form include the (odd-)occupancy processes in (1.1) and hence we refer to $S_{n}$ as weighted occupancy processes. Throughout, $a_{0}=0$. For the weights $(a_{j})_{j\in\mathbb{N}}$ we impose $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ for some $\beta\in[0,\alpha^{2}/2)$. When $\beta=0$ this implies that the weights are bounded, including the occupancy and odd-occupancy processes investigated earlier. Our main result is a functional central limit theorem for the counting process $(S_{\left\lfloor nt\right\rfloor})_{t\in[0,1]}$ appropriately normalized, and we identify a new family of Gaussian processes in the limit. Our results can be viewed as an extension of the functional central limit theorems in [15] (more precisely, Theorem 2.3 therein for the quenched version of the randomized model with fixed randomization $\varepsilon_{j}\equiv 1$, corresponding exactly the two cases in (1.1) here; see Remark 2.2 for more details). While we follow the standard Poissonization approach, it turned out that the analysis becomes much more involved even in the case where $(a_{j})_{j\in\mathbb{N}}$ is a bounded sequence. The technical challenges came from (a) proving the tightness for the Poissonized process where, due the fact that now $a_{j}$ may be unbounded (when $\beta>0$) we no longer have simple moment estimates, and (b) carrying out of the de-Possonization step, where we use a refined analysis on discrete local central limit theorems for Poisson and binomial distributions.
@@ 1.1. Main results
In words, $K_{n,\ell}$ is the number of balls in the urn labeled by $\ell$ by round $n$ and $D_{n,j}$ is the number of urns with exactly $j$ balls by round $n$. The key here is the decay rate of the frequencies. Let $10<\alpha<$ 1. The following assumption was introduced first by Darling, [12, Section 4] and Karlin, [25] about the same time (Darling only investigated weak laws of large number and Karlin investigated central limit theorems)
$v(x):=\sum_{i=1}^{\infty}\delta_{1/p_{i}}([0,x])=x^{\alpha}L(x),\quad x\geq 0,$
@@
as $n\to\infty$ for some constant $\mathsf{C}_{0}>0$, then $v(x)\sim\mathsf{C}_{0}^{\alpha}x^{\alpha}$. Throughout we write $a_{n}\sim b_{n}$ if $\lim_{n\to\infty}a_{n}/b_{n}=1$. More generally, one can consider $p_{n}=n^{-1/\alpha}L(n)$ for some slowly varying function $L$ at infinity. However we do not pursue this generality, as for our main applications on random matrix, $L(n)$ converges to a constant almost surely, and hence (1.2) suffices. The proofs are already quite involved without considering a more general $L$.
@@
    It is not clear to us when $a_{j}$ is unbounded whether the functional central limit theorem can be established without any further conditions beside $|a_{j}|\leq Cj^{\beta}$. It is known that when $a_{j}$ changes sign, the regularity of the sample paths can be quite poor. For a closely related issue, see [14, Remark 2].
@@
$\sum_{j=1}^{\infty}a_{j}\mathcal{Z}_{\alpha,j}(t)=\int_{\mathbb{R}_{+}\times\Omega_{+}}a_{j}({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}^{\prime}(N^{\prime}(rt)=j))\mathcal{M}_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),t\geq 0,$
@@
We mention the recent paper by Iksanov and Kotelnikova, [24] (see also [23, Proposition 4.1]), who investigated Karlin model with $\alpha=0$ (this is the case where $p_{j}$ is a slowly varying function at infinity and tending to zero), and obtained new Gaussian processes as the scaling limits for various statistics. In short, the case $\alpha=0$ of the Karlin model behaves drastically different from the case $\alpha\in(0,1)$ investigated here, yet it appears to have received much less attention. The Karlin model with $\alpha=0$ is closely related to the Bernoulli sieves [19, 22]. See also [24, 23] for related references.
@@ 1.2. An application in random matrix theory
The key connection to random permutation matrices is that, thanks to Kingman’s representation theorem [26], the occupancy counts $(D_{n,1},\dots,D_{n,n})$ of an infinite urn scheme have the same joint law of the cycle counts $(C_{n,1},\dots,C_{n,n})$ of the corresponding exchangeable random partition from a Chinese restaurant process (where $C_{n,j}$ is the number of the $j$-cycles of the partition of $[n]$), if the rate $(p_{j})_{j\in\mathbb{N}}$ is appropriately chosen (strictly speaking, this statement involves regular conditional probability; see [17] for more details). Since for a permutation the cycle structure determines completely the eigenvalues of the corresponding matrix, many statistics of interest can be expressed as weighted sum of cycle counts, which have the same law as the weighted occupancy counts investigated here. This connection allows us to study random permutation matrices by studying classical questions on infinite urn schemes. This idea was first due to Diaconis and Shahshahani, [13] and has been applied extensively to random permutation matrices either uniform or those induced by Ewens measures on symmetric groups with parameter $\theta>0$, corresponding to the Chinese restaurant processes with $(0,\theta)$-seating (see e.g. [1, 4, 6, 32, 5]). At the same time, the Karlin model with parameter $\alpha\in(0,1)$ can be associated to Chinese restaurant processes with $(\alpha,\theta)$-seating ($\theta>-\alpha$), and this indicates again the importance of Karlin model from another aspect. Surprisingly, we have found very few results on random permutations induced by $(\alpha,\theta)$-seating with $\alpha\in(0,1)$, and the connection to the infinite urn model has not been fully exploited yet.
Here we mention briefly an application in our accompanying paper [17], just to provide a flavor of the results to be established. We skip most the background including the definition of Chinese restaurant processes, as more details and other applications can be found therein.
@@
The random variable $\mathsf{s}_{\alpha,\theta}$ referred to as the $\alpha$-diversity, is strictly positive and is measurable with respect to
@@ Acknowledgements
Y.W. was grateful to Alexander Iksanov for several illuminating discussions. J.G. and Y.W. were partially supported by Army Research Office, US (W911NF-20-1-0139). Y.W. was also partially supported by Simons Foundation (MP-TSM-00002359), and a Taft Center Fellowship (2024–2025) from Taft Research Center at University of Cincinnati.
@@ 2. The limit process
then $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ forms a stochastic process with a version that is $\gamma$-Hölder-continuous with $\gamma\in(0,\alpha/2)$.
@@
Here and below, ${\mathbb{E}}^{\prime}$ is the expectation with respect to $\mathbb{P}^{\prime}$. The notation is solely to indicate that they have nothing to do with the probability space where the random measure $\mathcal{M}_{\alpha}$ is defined and for convenience from now on, when working with ${\mathbb{E}}^{\prime},\mathbb{P}^{\prime}$ and $N^{\prime}$ without the involvement of $\mathcal{M}_{\alpha}$ we may simply write ${\mathbb{E}},\mathbb{P}$ and $N$ accordingly.
Recall that $N^{\prime}(rs)\leq N^{\prime}(rt)$. The first integral above equals
@@
From stochastic-integral representations, in principle, one can obtain explicit formula of the covariance function following (2.4). However, most of the time the infinite series in the formulae cannot be simplified and are not directly useful. We mention two noteworthy exceptions, both from [15, Theorem 2.3].
@@
    In this case, $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ is known as the bi-fractional Brownian motion with parameters $(1/2,\alpha)$. See [15, Remark 2.4] for more discussions.
@@
and in particular it satisfies $z^{\alpha}=1-{\mathbb{E}}(1-z)^{Q_{\alpha}}$ for all $|z|<1$ (this is essentially Newton’s generalized binomial theorem). Then, we have
$\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\Gamma(k-\alpha)}{\Gamma(k+1)}=\Gamma(\alpha-1)t^{\alpha}{\mathbb{E}}\left(1-\left(1-\frac{s}{t}\right)^{Q_{\alpha}}\right)=\Gamma(\alpha-1)s^{\alpha},$
@@
$\alpha(s+t)^{\alpha}\sum_{k=2}^{\infty}\frac{F_{\vec{a},s,t}^{(2)}(k)\Gamma(k-\alpha)}{\Gamma(k+1)}$ $=\Gamma(\alpha-1)(s+t)^{\alpha}{\mathbb{E}}\left(1-\left(\frac{s}{s+t}\right)^{Q_{\alpha}}-\left(\frac{t}{s+t}\right)^{Q_{\alpha}}\right)$
$=\Gamma(\alpha-1)\left(s^{\alpha}+t^{\alpha}-(s+t)^{\alpha}\right).$
@@ 3. Limit theorem for the Poissonized model
We shall prove the following theorem, which is the Poissonized version of Theorem 1.1. Recall that we assume $p_{j}\sim\mathsf{C}_{0}j^{-1/\alpha}$ as $j\to\infty$, which we do not repeat in the sequel, and we take $\sigma_{n}^{2}=v(n)\sim\mathsf{C}_{0}^{\alpha}n^{\alpha}$. See the discussions around (1.2).
@@
To prove Theorem 3.1, we first prove the convergence of finite-dimensional distributions in Section 3.1. Note that in this step we assume only the weaker condition $|a_{j}|\leq Cj^{\beta}$ for $\beta\in[0,\alpha/2)$. Next, in Section 3.2 we establish the tightness, and it is here that we impose the stronger condition $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ with $\beta\in[0,\alpha^{2}/2)$.
@@ 3.1. Finite-dimensional convergence for the Poissonized model
We proceed by the following approximation. For every fixed $\epsilon>0$, write
@@
$\widetilde{W}^{(\epsilon)}_{n}(t):=\sum_{\ell=\ell_{n}}^{\infty}\left({a}_{N_{\ell}(nt)}-{\mathbb{E}}a_{N_{\ell}(nt)}\right)\quad\mbox{ and }\quad\widetilde{T}^{(\epsilon)}_{n}(t):=\sum_{\ell=1}^{\ell_{n}-1}\left({a}_{N_{\ell}(nt)}-{\mathbb{E}}a_{N_{\ell}(nt)}\right),$
@@
(3.2) $\sum_{\ell=\ell_{n}}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}\sim\sigma_{n}^{2}\alpha t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\gamma(j-\alpha,t\epsilon^{-{1}/{\alpha}})}{j!},$
@@
(3.3) $\sum_{\ell=1}^{\ell_{n}-1}{\mathbb{E}}b_{N_{\ell}(nt)}\leq C\sigma_{n}^{2}t^{\alpha^{\prime}}\epsilon^{1-\alpha^{\prime}/\alpha}$
@@
$\sum_{\ell=1}^{\infty}{\mathbb{E}}b_{N_{\ell}(nt)}\leq C\sigma_{n}^{2}t^{\alpha^{\prime}}.$
@@
Using the substitution $z=nt/x$ and then integration by parts we have
$\sum_{\ell=\ell_{n}}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}$ $=\int_{1/p_{\ell_{n}}}^{\infty}\sum_{j=1}^{\infty}\frac{b_{j}e^{-{nt}/{x}}}{j!}\left(\frac{nt}{x}\right)^{j}v({\rm d}x)=-\int_{0}^{ntp_{\ell_{n}}}\sum_{j=1}^{\infty}\frac{b_{j}}{j!}z^{j}e^{-z}v\left({\rm d}\frac{nt}{z}\right)$
@@
We claim that
@@
then for some $\delta^{\prime}>0$ for $n$ large enough such that $ntp_{\ell_{n}}<t\epsilon^{-1/\alpha}+\delta^{\prime}$, and another application of Potter’s theorem on $v(nt/z)/v(nt)$ we have for $\delta>0$ and some constant $C>0$ depending on $\delta$,
@@
$\lim_{n\rightarrow\infty}\frac{1}{\sigma_{n}^{2}}\sum_{\ell=\ell_{n}}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}$ $=\lim_{n\rightarrow\infty}\left(\frac{A_{n}}{\sigma_{n}^{2}}-\frac{B_{n}}{\sigma_{n}^{2}}\right)=t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\alpha\gamma(j-\alpha,t\epsilon^{-1/\alpha})}{j!}+R_{\epsilon}-R_{\epsilon}$
@@
$\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}b_{N_{\ell}(nt)}$ $\leq C\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}N_{\ell}(nt)^{\alpha^{\prime}}\leq C\sum_{\ell=1}^{\ell_{n}}({\mathbb{E}}N_{\ell}(nt))^{\alpha^{\prime}}=C\sum_{\ell=1}^{\ell_{n}}(p_{\ell}nt)^{\alpha^{\prime}}$
@@
By the Cramér–Wold device it suffices to show that, for distinct $t_{1}\dots,t_{d}\in[0,1]$ and any $c_{1},\dots,c_{d}\in\mathbb{R}$,
@@
We first compute the convergence of the asymptotic variance. Since
$\mathop{\rm Var}\left(\sum_{i=1}^{d}c_{i}\widetilde{W}^{(\epsilon)}_{n}(t_{i})\right)$ $=\sum_{i,j=1}^{d}c_{i}c_{j}\mathop{\rm Cov}\left(\widetilde{W}^{(\epsilon)}_{n}(t_{i}),\widetilde{W}^{(\epsilon)}_{n}(t_{j})\right).$
@@
The calculation is essentially the same as in Proposition 2.1, replacing $\int_{0}^{\infty}$ by $\int_{0}^{\epsilon^{-1/\alpha}}$. We omit the details.
Next, we compute the asymptotic covariance. Write
${\rm{Cov}}\left(\widetilde{W}_{n}^{(\epsilon)}(s),\widetilde{W}_{n}^{(\epsilon)}(t)\right)=\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)-\sum_{\ell=\ell_{n}}^{\infty}{\mathbb{E}}a_{N_{\ell}(ns)}{\mathbb{E}}a_{N_{\ell}(nt)}.$

### v3

@@ 1. Introduction
The infinite urn scheme has been a classical model in probability theory, dating back at least to the 1960s in the works of Bahadur, [3], Darling, [14] and Karlin, [27]. In particular, Darling and Karlin first considered frequencies decaying polynomially, that is, the sampling frequencies $(p_{j})_{j\in\mathbb{N}}$ of the infinite urn model satisfy $p_{j}=L(j)j^{-1/\alpha}$ for some slowly varying function $L$ at infinity and some parameter $\alpha\in(0,1]$. They also examined the case when $p_{j}$ decays faster than any polynomial rates (for a precise definition, see (1.2) below with $\alpha=0$). For these cases, Karlin established central limit theorems for various statistics, and therefore the model with such frequencies has been sometimes referred to as the Karlin model (although his seminal paper investigated other frequencies as well).
Among many statistics of interest, Karlin established central limit theorems for the so-called occupancy and odd-occupancy processes. If we let $D_{n,j}$ denote the number of urns with exactly $j$ balls after the first $n$ rounds of sampling, then these are the counting processes
@@
In the current paper, we establish a functional central limit theorem for the weighted sum of occupancy counts $(D_{n,j})_{j\in\mathbb{N}}$ with prescribed weights $\vec{a}=(a_{j})_{j\in\mathbb{N}}$ defined as
@@
The statistics in this form include the (odd-)occupancy processes in (1.1) and hence we refer to $S_{n}$ as weighted occupancy processes. Throughout, $a_{0}=0$. In particular, with $a_{j}=1$ and $a_{j}={{\bf 1}}_{\left\{j~{}\rm~{}is~{}odd\right\}}$, $S_{n,\vec{a}}$ becomes the two cases in (1.1) considered by Karlin respectively; see also [17]. Another extensively studied case is $a_{j}={{\bf 1}}_{\left\{j\geq j_{0}\right\}}$, and $S_{n,\vec{a}}=\sum_{j\geq j_{0}}D_{n,j}$ is the counting process of all urns with at least $j_{0}$ balls. This case was again investigated already by Karlin; see also [12]. As an immediate consequence one obtains limit theorems for $D_{n,j_{0}}$ by simply writing $D_{n,j_{0}}=\sum_{j\geq j_{0}}D_{n,j}-\sum_{j\geq j_{0}+1}D_{n,j}$ and a continuous mapping argument. We are not aware of any other examples of $S_{n,\vec{a}}$ investigated in the literature.
For the general weights $(a_{j})_{j\in\mathbb{N}}$ we impose $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ for some $\beta\in[0,\alpha^{2}/2)$ and all $i,j\in\mathbb{N}\cup\{0\}$. When $\beta=0$ this implies that the weights are bounded, including the occupancy and odd-occupancy processes investigated earlier. Our main result is a functional central limit theorem for the counting process $(S_{\left\lfloor nt\right\rfloor})_{t\in[0,1]}$ appropriately normalized, and we identify a new family of Gaussian processes in the limit. Our results can be viewed as an extension of the functional central limit theorems in [17] (more precisely, Theorem 2.3 therein for the quenched version of the randomized model with fixed randomization $\varepsilon_{j}\equiv 1$, corresponding exactly the two cases in (1.1) here; see Remark 2.2 for more details). While we follow the standard Poissonization approach, it turned out that the analysis becomes much more involved even in the case where $(a_{j})_{j\in\mathbb{N}}$ is a bounded sequence. The technical challenges came from (a) proving the tightness for the Poissonized process where, due the fact that now $a_{j}$ may be unbounded (when $\beta>0$) we no longer have simple moment estimates, and (b) carrying out of the de-Possonization step, where we use a refined analysis on discrete local central limit theorems for Poisson and binomial distributions.
@@ 1.1. Main results
In words, $K_{n,\ell}$ is the number of balls in the urn labeled by $\ell$ by round $n$ and $D_{n,j}$ is the number of urns with exactly $j$ balls by round $n$. The key here is the decay rate of the frequencies. Let $10<\alpha<$ 1. The following assumption was introduced first by Darling, [14, Section 4] and Karlin, [27] about the same time (Darling only investigated weak laws of large number and Karlin investigated central limit theorems)
(1.2) $v(x):=\sum_{i=1}^{\infty}\delta_{1/p_{i}}([0,x])=x^{\alpha}L(x),\quad x\geq 0,$
@@
as $n\to\infty$ for some constant $\mathsf{C}_{0}>0$, then $v(x)\sim\mathsf{C}_{0}^{\alpha}x^{\alpha}$ as $x\rightarrow\infty$. Throughout we write $a_{n}\sim b_{n}$ if $\lim_{n\to\infty}a_{n}/b_{n}=1$. More generally, one can consider $p_{n}=n^{-1/\alpha}L(n)$ for some slowly varying function $L$ at infinity. However we do not pursue this generality, as for our main applications on random matrices, $L(n)$ converges to a constant almost surely, and hence (1.3) suffices. The proofs are already quite involved without considering a more general $L$.
@@
    It is not clear to us when $a_{j}$ is unbounded whether the functional central limit theorem can be established without any further conditions beside $|a_{j}|\leq Cj^{\beta}$. It is known that when $a_{j}$ changes sign, the regularity of the sample paths can be quite poor. For a closely related issue, see [16, Remark 2].
@@
$\sum_{j=1}^{\infty}a_{j}\mathcal{Z}_{\alpha,j}(t)=\int_{\mathbb{R}_{+}\times\Omega_{+}}\sum_{j=1}^{\infty}a_{j}({{\bf 1}}_{\left\{N^{\prime}(rt)=j\right\}}-\mathbb{P}^{\prime}(N^{\prime}(rt)=j))\mathcal{M}_{\alpha}({\rm d}r,{\rm d}\omega^{\prime}),t\geq 0,$
@@
We mention the recent paper by Iksanov and Kotelnikova, [26] (see also [25, Proposition 4.1]), who investigated Karlin model with $\alpha=0$ in (1.2) (note that in this case (1.3) does not make sense; but instead $p_{j}$ decays faster than any polynomial rates), and obtained new Gaussian processes as the scaling limits for various statistics. In short, the case $\alpha=0$ of the Karlin model behaves drastically different from the case $\alpha\in(0,1)$ investigated here, yet it appears to have received much less attention. The Karlin model with $\alpha=0$ is closely related to the Bernoulli sieves [21, 24]. See also [26, 25] for related references.
@@ 1.2. An application in random matrix theory
The key connection to random permutation matrices is that, thanks to Kingman’s representation theorem [28], the occupancy counts $(D_{n,1},\dots,D_{n,n})$ of an infinite urn scheme have the same joint law of the cycle counts $(C_{n,1},\dots,C_{n,n})$ of the corresponding exchangeable random partition from a Chinese restaurant process (where $C_{n,j}$ is the number of the $j$-cycles of the partition of $[n]=\{1,2,...,n\}$), if the rate $(p_{j})_{j\in\mathbb{N}}$ is appropriately chosen (strictly speaking, this statement involves regular conditional probability; see [19] for more details). Since for a permutation the cycle structure determines completely the eigenvalues of the corresponding matrix, many statistics of interest can be expressed as weighted sum of cycle counts, which have the same law as the weighted occupancy counts investigated here. This connection allows us to study random permutation matrices by investigating classical questions on infinite urn schemes. This idea was first due to Diaconis and Shahshahani, [15] and has been applied extensively to random permutation matrices either uniform or those induced by Ewens measures on symmetric groups with parameter $\theta>0$, corresponding to the Chinese restaurant processes with $(0,\theta)$-seating (see e.g. [1, 4, 6, 35, 5]). At the same time, the Karlin model with parameter $\alpha\in(0,1)$ can be associated to Chinese restaurant processes with $(\alpha,\theta)$-seating ($\theta>-\alpha$), and this indicates again the importance of Karlin model from another aspect. Surprisingly, we have found very few results on random permutations induced by $(\alpha,\theta)$-seating with $\alpha\in(0,1)$, and the connection to the infinite urn model has not been fully exploited yet.
Here we mention briefly an application in our accompanying paper [19], just to provide a flavor of the results to be established. We skip most of the background including the definition of Chinese restaurant processes, as more details and other applications can be found therein.
@@
Here and throughout, $\Gamma$ is the gamma function. The random variable $\mathsf{s}_{\alpha,\theta}$ referred to as the $\alpha$-diversity, is strictly positive and is measurable with respect to
@@ Acknowledgements
The authors would like to thank two anonymous referees for very careful reading of and helpful comments on the manuscript. Y.W. was grateful to Alexander Iksanov for several illuminating discussions. J.G. and Y.W. were partially supported by Army Research Office, US (W911NF-20-1-0139). Y.W. was also partially supported by Simons Foundation (MP-TSM-00002359), and a Taft Center Fellowship (2024–2025) from Taft Research Center at University of Cincinnati.
@@ 2. The limit process
then $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ forms a stochastic process with a version that is $\gamma$-Hölder-continuous for every $\gamma\in(0,\alpha/2)$.
@@
Here and below, ${\mathbb{E}}^{\prime}$ is the expectation with respect to $\mathbb{P}^{\prime}$. The notation is solely to indicate that they have nothing to do with the probability space where the random measure $\mathcal{M}_{\alpha}$ is defined and for convenience from now on, when working with ${\mathbb{E}}^{\prime},\mathbb{P}^{\prime}$ and $N^{\prime}$ without the involvement of $\mathcal{M}_{\alpha}$ we may simply write ${\mathbb{E}},\mathbb{P}$ and $N$ accordingly. Recall that $N^{\prime}(rs)\leq N^{\prime}(rt)$ a.s. since $s<t$. The first integral above equals
@@
From stochastic-integral representations, we obtained an explicit formula of the covariance function (2.2). However, most of the time the infinite series in the formulae cannot be simplified and are not directly useful. We mention two noteworthy exceptions, both from [17, Theorem 2.3].
@@
    In this case, $(\mathcal{Z}_{\alpha}(t))_{t\geq 0}$ is known as the bi-fractional Brownian motion with parameters $(1/2,\alpha)$. See [17, Remark 2.4] for more discussions.
@@
and in particular it satisfies $z^{\alpha}=1-{\mathbb{E}}(1-z)^{Q_{\alpha}}$ for all $z\in[0,1)$ (this is essentially Newton’s generalized binomial theorem). Then, we have
$\alpha t^{\alpha}\sum_{k=1}^{\infty}\frac{F_{\vec{a},s,t}^{(1)}(k)\Gamma(k-\alpha)}{\Gamma(k+1)}=\Gamma(1-\alpha)t^{\alpha}{\mathbb{E}}\left(1-\left(1-\frac{s}{t}\right)^{Q_{\alpha}}\right)=\Gamma(1-\alpha)s^{\alpha},$
@@
$\alpha(s+t)^{\alpha}\sum_{k=2}^{\infty}\frac{F_{\vec{a},s,t}^{(2)}(k)\Gamma(k-\alpha)}{\Gamma(k+1)}$ $=\Gamma(1-\alpha)(s+t)^{\alpha}{\mathbb{E}}\left(1-\left(\frac{t}{s+t}\right)^{Q_{\alpha}}-\left(\frac{s}{s+t}\right)^{Q_{\alpha}}\right)$
$=\Gamma(1-\alpha)\left(s^{\alpha}+t^{\alpha}-(s+t)^{\alpha}\right).$
@@ 3. Limit theorem for the Poissonized model
We shall prove the following theorem, which is the Poissonized version of Theorem 1.1. Recall that we assume $p_{j}\sim\mathsf{C}_{0}j^{-1/\alpha}$ as $j\to\infty$, which we do not repeat in the sequel, and we shall denote $\sigma_{n}^{2}:=v(n)\sim\mathsf{C}_{0}^{\alpha}n^{\alpha}$. See the discussions around (1.3).
@@
To establish Theorem 3.1, we first prove the convergence of finite-dimensional distributions in Section 3.1. Note that in this step we assume only the weaker condition $|a_{j}|\leq Cj^{\beta}$ for $\beta\in[0,\alpha/2)$. Next, in Section 3.2 we establish the tightness, and it is here that we impose the stronger condition $|a_{i+j}-a_{i}|\leq Cj^{\beta}$ with $\beta\in[0,\alpha^{2}/2)$.
@@ 3.1. Finite-dimensional convergence for the Poissonized model
We proceed by the following approximation. Throughout this subsection without loss of generality assume that $(p_{\ell})_{\ell\in\mathbb{N}}$ is non-increasing. For every fixed $\epsilon>0$, write
@@
$\widetilde{W}^{(\epsilon)}_{n}(t):=\sum_{\ell=\ell_{n}+1}^{\infty}\left({a}_{N_{\ell}(nt)}-{\mathbb{E}}a_{N_{\ell}(nt)}\right)\quad\mbox{ and }\quad\widetilde{T}^{(\epsilon)}_{n}(t):=\sum_{\ell=1}^{\ell_{n}}\left({a}_{N_{\ell}(nt)}-{\mathbb{E}}a_{N_{\ell}(nt)}\right),$
@@
(3.2) $\sum_{\ell=\ell_{n}+1}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}\sim\sigma_{n}^{2}\alpha t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\gamma(j-\alpha,t\epsilon^{-{1}/{\alpha}})}{j!},$
@@
(3.3) $\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}|b_{N_{\ell}(nt)}|\leq C\sigma_{n}^{2}t^{\alpha^{\prime}}\epsilon^{1-\alpha^{\prime}/\alpha}$
@@
$\sum_{\ell=1}^{\infty}{\mathbb{E}}|b_{N_{\ell}(nt)}|\leq C\sigma_{n}^{2}t^{\alpha^{\prime}}.$
@@
We start by writing
$\sum_{\ell=\ell_{n}+1}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}$ $=\int_{1/p_{\ell_{n}}}^{\infty}\sum_{j=1}^{\infty}\frac{b_{j}e^{-{nt}/{x}}}{j!}\left(\frac{nt}{x}\right)^{j}v({\rm d}x)$
$=-\int_{1/p_{\ell_{n}}}^{\infty}\sum_{j=1}^{\infty}\frac{b_{j}e^{-nt/x}}{j!}\left(\frac{nt}{x^{2}}-\frac{j}{x}\right)\left(\frac{nt}{x}\right)^{j}v(x){\rm d}x-v(p_{\ell_{n}}^{-1})\sum_{j=1}^{\infty}\frac{b_{j}e^{-ntp_{\ell_{n}}}(ntp_{\ell_{n}})^{j}}{j!}$
@@
Here, the first equality follows from the definition of Riemann–Stieltjes integral and also the assumption that $(p_{\ell})_{\ell\in\mathbb{N}}$ is non-increasing, the second follows from integration by parts, and the third follows from a change of variables. We claim that
@@
then for some $\delta^{\prime}>0$ for $n$ large enough such that $ntp_{\ell_{n}}<t\epsilon^{-1/\alpha}+\delta^{\prime}$, and an application of Potter’s theorem (see Theorem 1.5.3 in [8]) on $v(nt/z)/v(nt)$ we have for $\delta\in(0,1-\alpha)$ and some constant $C>0$ depending on $\delta$,
@@
$\lim_{n\rightarrow\infty}\frac{1}{\sigma_{n}^{2}}\sum_{\ell=\ell_{n}+1}^{\infty}\mathbb{E}b_{N_{\ell}(nt)}$ $=\lim_{n\rightarrow\infty}\left(\frac{A_{n}}{\sigma_{n}^{2}}-\frac{B_{n}}{\sigma_{n}^{2}}\right)=t^{\alpha}\sum_{j=1}^{\infty}\frac{b_{j}\alpha\gamma(j-\alpha,t\epsilon^{-1/\alpha})}{j!}+R_{\epsilon}-R_{\epsilon}$
@@
$\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}|b_{N_{\ell}(nt)}|$ $\leq C\sum_{\ell=1}^{\ell_{n}}{\mathbb{E}}N_{\ell}(nt)^{\alpha^{\prime}}\leq C\sum_{\ell=1}^{\ell_{n}}({\mathbb{E}}N_{\ell}(nt))^{\alpha^{\prime}}=C\sum_{\ell=1}^{\ell_{n}}(p_{\ell}nt)^{\alpha^{\prime}}$
@@
By the Cramér–Wold device it suffices to show that, for any $d\in\mathbb{N}$, distinct $t_{1}\dots,t_{d}\in[0,1]$ and any $c_{1},\dots,c_{d}\in\mathbb{R}$,
@@
Note that
$\sum_{i=1}^{d}c_{i}\widetilde{W}_{n}^{(\epsilon)}(t_{i})=\sum_{\ell=\ell_{n}+1}^{\infty}\sum_{i=1}^{d}c_{i}\left(a_{N_{\ell}(nt_{i})}-{\mathbb{E}}a_{N_{\ell}(nt_{i})}\right)$
is a summation of independent random variables, and hence we can apply the central limit theorem for an infinite triangular array of independent random variables satisfying a Lyapunov-type condition (3.7) below ([10, Theorem 8.4.1]).
We first show the convergence of variance. Since
$\mathop{\rm Var}\left(\sum_{i=1}^{d}c_{i}\widetilde{W}^{(\epsilon)}_{n}(t_{i})\right)$ $=\sum_{i,j=1}^{d}c_{i}c_{j}\mathop{\rm Cov}\left(\widetilde{W}^{(\epsilon)}_{n}(t_{i}),\widetilde{W}^{(\epsilon)}_{n}(t_{j})\right),$
@@
The calculation is essentially the same as in Proposition 2.1, replacing $\int_{0}^{\infty}$ by $\int_{0}^{\epsilon^{-1/\alpha}}$. We omit the details. Next, we compute the limit covariance. Write
${\rm{Cov}}\left(\widetilde{W}_{n}^{(\epsilon)}(s),\widetilde{W}_{n}^{(\epsilon)}(t)\right)=\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}\left(a_{N_{\ell}(ns)}a_{N_{\ell}(nt)}\right)-\sum_{\ell=\ell_{n}+1}^{\infty}{\mathbb{E}}a_{N_{\ell}(ns)}{\mathbb{E}}a_{N_{\ell}(nt)}.$